# Efficient Retrieval with Learned Similarities

### v2

@@ Abstract
Retrieval plays a fundamental role in recommendation systems, search, and natural language processing by efficiently finding relevant items from a large corpus given a query. Dot products have been widely used as the similarity function in such retrieval tasks, thanks to Maximum Inner Product Search (MIPS) that enabled efficient retrieval based on dot products. However, state-of-the-art retrieval algorithms have migrated to learned similarities. Such algorithms vary in form; the queries can be represented with multiple embeddings, complex neural networks can be deployed, the item ids can be decoded directly from queries using beam search, and multiple approaches can be combined in hybrid solutions. Unfortunately, we lack efficient solutions for retrieval in these state-of-the-art setups. Our work investigates techniques for approximate nearest neighbor search with learned similarity functions. We first prove that Mixture-of-Logits (MoL) is a universal approximator, and can express all learned similarity functions. We next propose techniques to retrieve the approximate top $K$ results using MoL with a tight bound. We finally compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
1 Introduction
@@
Retrieval requires efficient storing, indexing, and querying relevant candidate items represented by high-dimensional vectors. Retrieval is widely used as the initial preprocessing stage for applications such as recommendations and natural language processing that operate over corpus with up to billions of items [5, 9, 26, 25, 2]. In many concrete use cases, such as vector databases [19], the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
@@
Despite the popularity of dot products and numerous work done to improve their efficiency [18, 35, 28, 4], state-of-the-art retrieval algorithms have long moved to various learned similarity functions. Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values [26, 22]. As another example, Probabilistic Label Trees (PLTs) [16] and Tree-based Deep Models (TDMs) [45, 46] map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf. More recent work on generative retrieval in recommendations and natural language processing directly map the query to the item ids in sequence-to-sequence or decoder-only setups [38, 40]. Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products [8]. Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms [14, 41, 34, 42].
@@
Supporting efficient retrieval with these diverse learned similarities is challenging. Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps [45, 42], while recommendation and NLP deployments commonly need to handle billions of items [6, 26, 3] with low end-to-end latency. When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force [14] or heuristics [42]. What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface makes it even more difficult for designing a general solution for efficient retrieval.
@@
Taking a step back, our key insight is that learned similarity approaches are but different ways to increase expressiveness of the retrieval stage. Formally, for a query $q$ and an item $x$, the expressiveness of the similarity function boils down to deriving alternative parameterizations of $p(x|q)$ matrices, with full rank matrices being the most expressive among them. Dot products, on the other hand, induces a low-rank bottleneck due to the dimensionality of the embedding, i.e., $\ln p(x|q)\propto\langle f(q),g(x)\rangle$ ($f(q),g(x)\in\mathbb{R}^{d}$). This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases [19, 4], and overfitting issues that come with larger embedding dimensions due to the need to co-train query- and item- encoders from data [5, 26, 8, 43].
@@
This insight enables us to support efficient retrieval with learned similarity functions by approximating them with MoL. To the best of our knowledge, this is the first work that tackles the problem of efficient retrieval with universal learned similarities. We first show that Mixture-of-Logits (MoL) is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions. Compared with prior work on learned similarities [14, 38] which often fail to outperform dot product baselines in large-scale settings [33, 36], our work lays theoretical foundations for MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 on corpus with hundreds of millions to billions of items [42, 3]. We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight bound. Our solution leverages existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS [18, 35, 19, 12]. We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms. Importantly, our approach with learned similarities efficiently utilizes modern accelerators due to MoL’s higher arithmetic intensity [42], which results in MIPS-level inference latency and throughput. This concludes both theoretical and practical justifications to migrate away from the broadly adopted MIPS solution in vector databases to Retrieval with Learned Similarities (RAILS) on GPUs.
2 Expressiveness of Mixture of Logits
In this section, we describe Mixture of Logits (MoL) and show that MoL is expressive enough to represent any learned similarity function. Table 1 summarizes the notations in this paper.
We first describe the learned similarity function with Mixture of Logits (MoL).
@@ 2 Expressiveness of Mixture of Logits
$P$ ($P_{q}$, $P_{x}$) MoL uses $P$ low-rank embeddings ("component-level embeddings") to represent $q$ and $x$. With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of embeddings for $q$ and $x$, respectively, and $P=P_{q}\times P_{x}$.
@@
$\langle f(q),g(x)\rangle$ the dot product similarity function; $\langle f(q),g(x)\rangle=g(x)^{T}f(q)$.
Table 1: Table of Notations.
@@
MoL [42, 3] assumes that the query $q$ and the item $x$ are already mapped to $P$ groups of low-rank embeddings (“component-level embeddings”), $f_{p}(q),g_{p}(x)\in\mathbb{R}^{d_{P}}$, where $f_{p}(q),g_{p}(x)$ are parameterized with some neural networks based on query and item features, respectively. MoL then calculates the similarity between the query $q$ and the item $x$ by applying adaptive gating weights, $\pi_{p}(q,x)\in[0,1]$, to the inner products of these $P$ pairs of low-rank embeddings, or $\langle f_{p}(q),g_{p}(x)\rangle$s. Note that prior work assumes that $\sum_{p}\pi_{p}(q,x)=1$ [42, 3], and parameterizes $\pi_{p}(q,x)$ utilizing some neural network with $f_{p}(q)$s, $g_{p}(x)$s, and $\langle f_{p}(q),g_{p}(x)\rangle$s as input features, but these details do not affect our analyses in this paper. Following [42]:
$\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$ (1)
To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators like GPUs, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embedding, where $P_{q}\times P_{x}=P$, and applying l2-norm to $f_{p}(q)$s and $g_{p}(x)$s:
$\phi(q,x)=\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}\pi_{p_{q},p_{x}}(q,x)\left\langle\frac{f_{p_{q}}(q)}{||f_{p_{q}}(q)||_{2}},\frac{g_{p_{x}}(x)}{||g_{p_{x}}(x)||_{2}}\right\rangle$ (2)
@@
Figure 1: Mixture-of-logits (MoL) learned similarity.
Now we show that any high rank matrix can be decomposed into a mixture of logits based on low rank matrices. Specifically, without loss of generality, we prove the following:
@@
Remark Here, we have shown that any high-rank matrix can be expressed as a mixture of logits of two low-rank matrices. Note that our decomposition is not intended to be used as a distillation of the original high-rank matrix. For instance, it is prohibitively expensive to populate the full matrix with a learned similarity function. In addition, the proof also does not indicate that having a sum of two mixture of logits is sufficient to train the embeddings and the learned similarity function. It is well-known that overparameterization is often necessary to enable efficient and performant training.
3 Retrieval Algorithms
In this section, we describe the problem of retrieving the top $K$ items with MoL and both exact and approximate algorithms for retrieval. Formally, we define the top $K$ retrieval problem as follows:

### v4

# Retrieval with Learned Similarities

@@ Abstract.
Retrieval plays a fundamental role in recommendation systems, search, and natural language processing (NLP) by efficiently finding relevant items from a large corpus given a query. Dot products have been widely used as the similarity function in such tasks, enabled by Maximum Inner Product Search (MIPS) algorithms for efficient retrieval. However, state-of-the-art retrieval algorithms have migrated to learned similarities. These advanced approaches encompass multiple query embeddings, complex neural networks, direct item ID decoding via beam search, and hybrid solutions. Unfortunately, we lack efficient solutions for retrieval in these state-of-the-art setups. Our work addresses this gap by investigating efficient retrieval techniques with expressive learned similarity functions. We establish Mixture-of-Logits (MoL) as a universal approximator of similarity functions, demonstrate that MoL’s expressiveness can be realized empirically to achieve superior performance on diverse retrieval scenarios, and propose techniques to retrieve the approximate top-$k$ results using MoL with tight error bounds. Through extensive experimentation, we show that MoL, enhanced by our proposed mutual information-based load balancing loss, sets new state-of-the-art results across heterogeneous scenarios, including sequential retrieval models in recommendation systems and finetuning language models for question answering; and our approximate top-$k$ algorithms outperform baselines by up to $\times6$ in latency while achieving >.$>.99$ recall rate compared to exact algorithms.  [^1] Our code and model checkpoints are available at https://github.com/bailuding/rails.
Nearest Neighbor Search, Learned Similarities, Top-K Retrieval, Vector Databases, Recommendation Systems, Question Answering
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714822isbn: 979-8-4007-1274-6/25/04ccs: Information systems Similarity measuresccs: Information systems Top-k retrieval in databasesccs: Information systems Learning to rankccs: Information systems Probabilistic retrieval modelsccs: Information systems Question answeringccs: Information systems Recommender systemsccs: Computing methodologies Natural language processingccs: Information systems Personalization
1. Introduction
@@
Retrieval requires efficient storing, indexing, and querying relevant candidate items represented by high-dimensional vectors. Retrieval is widely used as the initial preprocessing stage for internet applications such as recommendations, search, question answering, and natural language processing that operate over corpus with up to billions of items (Covington et al., 2016; Gillick et al., 2018; Li et al., 2019; Karpukhin et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022). In many concrete use cases, such as vector databases (Johnson et al., 2021), the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
@@
Despite the popularity of dot products and numerous work done to improve their efficiency (Jegou et al., 2011; Shrivastava and Li, 2014; Malkov and Yashunin, 2020; Chern et al., 2022), state-of-the-art retrieval algorithms have long moved to various learned similarity functions. Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values (Li et al., 2019; Khattab and Zaharia, 2020). As another example, Probabilistic Label Trees (PLTs) (Jasinska et al., 2016) and Tree-based Deep Models (TDMs) (Zhu et al., 2018; Zhuo et al., 2020) map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf. More recent work on generative retrieval directly map the query to the item ids in sequence-to-sequence or decoder-only setups (De Cao et al., 2021; Bevilacqua et al., 2022; Tay et al., 2022; Wang et al., 2022b; Sun et al., 2023). Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products (Gao et al., 2021). Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms (He et al., 2017; Yang et al., 2018; Santhanam et al., 2022; Zhai et al., 2023).
@@
Supporting efficient retrieval with these diverse learned similarities is challenging. Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps (Zhu et al., 2018), while recommendation and NLP deployments commonly need to handle billions of items (Eksombatchai et al., 2018; Li et al., 2019; Borisyuk et al., 2024) with a latency budget of tens of milliseconds. When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force (He et al., 2017) or heuristics (Zhai et al., 2023). While graph-based methods can be used to prune the search space (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019; Wang et al., 2022a; Ootomo et al., 2024), such methods tend to be much slower compared with MIPS algorithms leveraging quantization at high recall rates (Guo et al., 2020; ann, [n. d.]), and their performance can degrade when the similarity function is not a distance metric (Morozov and Babenko, 2018). What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface makes it even more difficult to design a general solution for efficient retrieval.
@@
Taking a step back, our key insight is that learned similarity approaches are but different ways to increase the expressiveness of the retrieval stage. Formally, for a query $q$ and an item $x$, the expressiveness of the similarity function boils down to deriving alternative parameterizations of $p(x|q)$ matrices, with full rank matrices being the most expressive among them. Dot products, on the other hand, induces a low-rank bottleneck due to the dimensionality of the embedding, i.e., $\ln p(x|q)\propto\langle f(q),g(x)\rangle$ ($f(q),g(x)\in\mathbb{R}^{d}$). This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023), and overfitting issues that come with larger embedding dimensions due to the common need to co-train or finetune query- and item- encoders from data (Covington et al., 2016; Li et al., 2019; Karpukhin et al., 2020; Gao et al., 2021; Ni et al., 2022a, b; Zhai et al., 2024).
@@
This insight enables us to support efficient retrieval with expressive learned similarity functions by approximating them with Mixture-of-Logits (MoL). To the best of our knowledge, this is the first work that tackles the problem of efficient retrieval with universal learned similarities, while setting new state-of-the-art results across heterogeneous scenarios. We first show that Mixture-of-Logits is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions (Section 2.1). Our work lays theoretical foundations for MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 on web-scale corpus with hundreds of millions to billions of items (Zhai et al., 2023; Borisyuk et al., 2024), and further enables MoL to be effectively applied across diverse retrieval scenarios, from large-scale recommendation systems to finetuning language models for question answering (Section 2.2). We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight error bound (Section 3). Our solution leverages the existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS (Jegou et al., 2011; Shrivastava and Li, 2014; Johnson et al., 2021; Guo et al., 2020). We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval and question answering tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to $\times6$ in latency, while achieving >.$>.99$ recall rate of exact algorithms (Section 4). Importantly, our approach with learned similarities efficiently utilizes modern accelerators due to MoL’s higher arithmetic intensity (Zhai et al., 2023), which results in MIPS-level inference latency and throughput. Overall, our work provides strong theoretical and practical justifications to migrate away from the broadly adopted MIPS solution in vector databases to Retrieval with Learned Similarities (RAILS) on GPUs.
2. Mixture of Logits
@@
$P$ ($P_{q}$, $P_{x}$) MoL uses $P$ pairs of low-rank embeddings ("component-level embeddings") to represent $q$ and $x$. With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of embeddings for $q$ and $x$, respectively; $P=P_{q}\times P_{x}$.
@@
$\langle f(q),g(x)\rangle$ the dot product similarity function: $g(x)^{T}f(q)$. $\langle f_{p}(q),g_{p}(x)\rangle$ denotes the dot product for the $p^{th}$ embedding pair.
Table 1. Table of Notations.
In this section, we describe Mixture of Logits (MoL), propose a load balancing loss to improve conditional computations in MoL, prove that MoL is expressive enough to represent any learned similarity function, and demonstrate how to apply MoL to diverse retrieval tasks. Table 1 summarizes the notations in this paper.
We first describe Mixture of Logits (MoL).
@@
MoL (Zhai et al., 2023) assumes that the query $q$ and the item $x$ are already mapped to $P$ pairs of low-rank embeddings (“component-level embeddings”), $f_{p}(q),g_{p}(x)\in\mathbb{R}^{d_{P}}$, where $f_{p}(q),g_{p}(x)$ are parameterized with some neural networks based on query and item features, respectively, and $d_{P}$ is the dimensionality of the low-rank embeddings. MoL then calculates the similarity between the query $q$ and the item $x$ by applying adaptive gating weights, $\pi_{p}(q,x)\in[0,1]$, to the inner products of these $P$ pairs of low-rank embeddings, or $\langle f_{p}(q),g_{p}(x)\rangle$s. Note that prior work assumes that $\sum_{p}\pi_{p}(q,x)=1$ (Zhai et al., 2023; Borisyuk et al., 2024), but this does not affect our analyses in this paper. Following (Zhai et al., 2023):
(1) $\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$
To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators like GPUs, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embeddings, where $P_{q}\times P_{x}=P$, and applying l2-norm to the embeddings:
(2) $\phi(q,x)=\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}\pi_{p_{q},p_{x}}(q,x)\left\langle\frac{f_{p_{q}}(q)}{||f_{p_{q}}(q)||_{2}},\frac{g_{p_{x}}(x)}{||g_{p_{x}}(x)||_{2}}\right\rangle$
@@
Mixture of Logits (MoL) with load balancing regularization loss.
We further observe $\pi_{p}(q,x)$ defines conditional computation to be performed over the $p$ low-rank embedding pairs, or $(f_{p}(q),g_{p}(x))$s. $\pi_{p}(q,x)$ should hence satisfy two conditions:
  • Globally, the $p$ low-rank embedding pairs, or $(f_{p}(q),g_{p}(x))$s, should receive a similar number of training examples even when $p$ is large and $\pi_{p}(q,x)$ is sparse, with load distributed evenly across the $p$ pairs. One way to do this is to maximize the entropy $H(p)$ over these embedding pairs.
  • The low-rank embedding pairs used to compute a particular $\phi(q,x)$ should be non-uniform and ideally sparse; e.g., it’s desirable to avoid the degenerate solution where $\pi_{p}(q,x)=\frac{1}{p}$. One way to do this is to minimize the conditional entropy $H(p|(q,x))$ of $p$ given (query, item) pairs.
Given these two desired conditions, we propose a mutual information-based regularization loss for load balancing, defined as
(3) $\mathcal{L}_{MI}=-H(p)+H(p|(q,x))$
with the overall training loss as
(4) $-\log\frac{\exp(\phi(q,x))}{\exp(\phi(q,x))+\sum_{x^{\prime}\in\mathbb{X}}\exp(\phi(q,x^{\prime}))}+\alpha\mathcal{L}_{MI}$
where the first part of Equation 4 is the sampled softmax loss used in (Zhai et al., 2023), and the second part adjusts the weight for the mutual information-based load balancing loss with a hyperparameter $\alpha$.
Figure 1. Mixture-of-logits (MoL) learned similarity.
2.1. Expressiveness of Mixture of Logits
Now we show that any high-rank matrix can be decomposed into a mixture of logits based on low-rank matrices, i.e., MoL is a universal approximator for all similarity functions. Without loss of generality, we prove the following:
@@ 2.1. Expressiveness of Mixture of Logits
Remark Here, we have shown that any high-rank matrix can be expressed as a mixture of logits of two low-rank matrices. Note that our decomposition is not intended to be used as a distillation of the original high-rank matrix. It is likely prohibitively expensive to populate the full matrix with a learned similarity function. In addition, our proof also does not indicate that having two mixture components is sufficient to train the embeddings and the learned similarity function. It is well-known that overparameterization is often necessary to enable efficient and performant training.
2.2. Applying MoL to Heterogeneous Use Cases
We now discuss how to apply MoL to retrieval tasks in different domains. Parameterization of the low-rank, component-level embeddings, or $f_{p}(q),g_{p}(x)\in\mathbb{R}^{d_{P}}$, plays an important role in realizing MoL’s theoretical expressiveness in practice, as suggested by prior work (Borisyuk et al., 2024). We discuss two scenarios on the opposite end of the spectrum, one with a large number of heterogeneous features – retrieval in large-scale recommendation systems, followed by another with a single homogeneous feature – finetuning language models for question answering and related NLP use cases, shown in Figure 2.
Retrieval in Large-scale Recommendation Systems.
Recommendation systems are characterized by the large number of heterogeneous features they use (Covington et al., 2016; Song et al., 2019; Zhai et al., 2024). This naturally enables some of those features to be utilized on the query- (user-) or on the item-side. For instance, embeddings can be constructed based on cluster ids on both the query-side and the item-side (Borisyuk et al., 2024). For common benchmark datasets, User ID-based one-hot embeddings (Koren et al., 2009) represent another possible $g_{p}(q)$ to use, which we evaluate in Section 4.
Finetuning Language Models for Question Answering.
In contrast, language models are characterized by their use of homogeneous semantic features, such as wordpieces and sentencepieces (Kudo and Richardson, 2018). We observe that MoL can be similarly adopted for those use cases. To obtain the $P_{X}$ item embeddings for MoL, we expand the tokenizer’s vocabulary with $P_{X}$ special aggregation tokens $X_{1},\ldots,X_{P_{X}}$, and append those $P_{X}$ tokens at the beginning of every tokenized sequence, $SP_{1},\ldots,SP_{N}$, as illustrated in Figure 2  [^1] Note that many question answering scenarios (Karpukhin et al., 2020; De Cao et al., 2021; Wang et al., 2022b; Ni et al., 2022b; Sun et al., 2023) utilize bidirectional language models for retrieval, like BERT (Devlin et al., 2019) or T5 (Raffel et al., 2023); for recent unidirectional language models, we can add $X_{1},\ldots,X_{P_{X}}$ to the end of the input sequence instead.. These $P_{X}$ special tokens play similar roles as the CLS token in BERT (Devlin et al., 2019), and during finetuning of the language model, are co-trained to aggregate different aspects of information as inputs for MoL. Additionally, we can design a learned pooling function to adapt pooling policy at an example-level (“Parameterized Pooling”) to improve model quality, which we discuss further in Appendix A.2.
Figure 2. Illustration of how to apply Mixture-of-logits (MoL) learned similarity to various retrieval scenarios, with a language model finetuning use case (characterized by a single homogeneous feature) shown on the left, and a recommendation use case (characterized by a large number of heterogeneous features) shown on the right. More details can be found in Appendix A.2.
3. Retrieval Algorithms
In this section, we describe the problem of retrieving the top $K$ items with MoL as well as exact and approximate retrieval algorithms. Formally, we define the top $K$ retrieval problem as follows: